An ammeter has a resistance of 0.05 ohms and reads 1ampere per scale division. What shunt is required so that it will read 10 ampere per scale division ?
1
Expert's answer
2020-06-22T11:09:22-0400
"I_s=I-I_a=10-1=9\ A"

"U=I_sR=I_ar"

"9R=(1)(0.05)"

"R=0.0056\Omega"